OpenStudy (anonymous):

PLEASE HELP ME Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 (14, 12) (12, 14) There are an infinite number of solutions. There is no solution.

TheSmartOne (thesmartone):

So do you know how to solve it?

OpenStudy (anonymous):

i got no solution

TheSmartOne (thesmartone):

It says substitution so we plug in y=4x-6 into 8x – 2y = 14 So we get 8x-2(4x-6)=14 8x-8x+12=14 12\(\ \neq\)14 So yes you are correct. It is no solution.

TheSmartOne (thesmartone):

If you get the question marks, just refresh your page. :D

OpenStudy (anonymous):

thanks

TheSmartOne (thesmartone):

No problem. Good job. You actually got the answer. I just had to check it for you. Keep it up! :D